OpenStudy (eboogie4):

Determine the circumference of a circle with a diameter of 13. Round your answer to the tenths place. a.132.7 units b.40.8 units c.530.9 units d.81.7 units

OpenStudy (mathstudent55):

Do you know the formula of the circumference of a circle?

OpenStudy (eboogie4):

ttps://www.explorelearning.com/index.cfm?method=cExtAccessSecure.dspResource&ResourceID=206&certificate=authorizer%3dE2020%26userid%3dGizmos4Test%26expires%3d2016%2f02%2f22+16%3a33%3a55%26hash%3dk5DTMiW8YC1zJG47o88GZw%253d%253d

OpenStudy (mathstudent55):

\(C = \pi d = 2 \pi r\)

OpenStudy (mathstudent55):

Use 3.14 for \(\pi\). d is the diameter.

OpenStudy (eboogie4):

@Adrianna.Gongora @Anaise @Bookworm14 @Daniellelovee

OpenStudy (shadowlegendx):

@mathstudent55 laid it out perfectly. Use the formula \[Circumference = \pi d\] \[\pi = 3.14 \] d = diameter

OpenStudy (shadowlegendx):

What is the diameter in this problem?

OpenStudy (shadowlegendx):

You want us to do the problem for you...?

OpenStudy (bookworm14):

TO find the circumference you must use 2 * pi * radius OR pi * diameter insert your diameter 13 * pi = what?

OpenStudy (bookworm14):

@Eboogie4 what is 13 * pi

OpenStudy (shadowlegendx):

This would be your circle: http://prntscr.com/a6r2ox But you should try work through the problem, and not relying upon an online tool.

OpenStudy (eboogie4):

40.82

OpenStudy (mathstudent55):

C = (pi)d = 3.14 * 13 = 40.82 You are correct.